Carla Vista Inc., is expected to grow at a rate of 19.000 percent for the next five years and then settle to a constant growth rate of 6.000 percent. The company recently paid a dividend of $2.35. The required rate of return is 16.000 percent

Find the present value of the dividends during the rapid-growth period if dividends grow at the same rate as the company. (Round dividends to 3 decimal places, e.g. 3.351. Round present value of dividends to 2 decimal places, e.g. 15.20.)
Answer #1
Year Dividend
1 $      2.80
2 $      3.33
3 $      3.96
4 $      4.71
5 $      5.61
PV $12.69

D1 = D0 x (1 + g) = 2.35 x (1 + 19%) = $2.80 and so on...

PV = D1 / (1 + r) + D2 / (1 + r)^2 + D3 / (1 + r)^3 + D4 / (1 + r)^4 + D5 / (1 + r)^5

= 2.80 / 1.16 + 3.33 / 1.16^2 + 3.96 / 1.16^3 + 4.71 / 1.16^4 + 5.61 / 1.16^5

= $12.69